# Quantum Quasinormal Mode Theory for Dissipative Nano-Optics and Magnetodielectric Cavity Quantum Electrodynamics

### v1

@@ Abstract
The unprecedented pace of evolution in nanoscale architectures for cavity quantum electrodynamics (cQED) has posed crucial challenges for theory, where the quantum dynamics arising from the non-perturbative dressing of matter by cavity electric and magnetic fields, as well as the fundamentally non-hermitian character of the system are to be treated without significant approximation. The lossy electromagnetic resonances of photonic, plasmonic or magnonic nanostructures are described as quasinormal modes (QNMs), whose properties and interactions with quantum emitters and spin qubits are central to the understanding of dissipative nano-optics and magnetodielectric cQED. Despite recent advancements toward a fully quantum framework for QNMs, a general and universally accepted approach to QNM quantization for arbitrary linear media remains elusive. In this work, we introduce a unified theoretical framework, based on macroscopic QED and complex coordinate transformations, that achieves QNM quantization for a wide class of spatially inhomogeneous, dissipative (with possible gain components) and dispersive, linear, magnetodielectric resonators. The complex coordinate transformations effectively convert the radiative losses into non-radiative material dissipation, and via a suitable transformation that reflects all the losses of the resonator, we define creation and annihilation operators that allow the construction of modal Fock states for the joint excitations of field-dressed matter. By directly addressing the intricacies of modal loss in a fully quantum theory of magnetodielectric cQED, our approach enables the exploration of modern, quantum nano-optical experiments utilizing dielectric, plasmonic, magnetic or hybrid cQED architectures, and paves the way towards a rigorous assessment of room-temperature quantum nanophotonic technologies without recourse to ad hoc quantization schemes.
@@ I Introduction
Cavity-mediated control and enhancement of light-matter interaction has become a pervasive paradigm in modern quantum science and technology. Already decades in maturity, the use of micro- and nanoscale, optical dielectric resonators to both harness and manipulate the photonic properties of semiconductor quantum dots has enabled the exploration of fundamental cavity quantum electrodynamics (cQED) in the solid state Vahala (2003); Khitrova et al. (2006); Kira and Koch (2011), as well as a host of device applications ranging from high-performance, non-classical light sources and spin-photon interfaces to ultrafast, all-optical switching for integrated photonic quantum computation and networks Hepp et al. (2019); Arakawa and Holmes (2020); Heindel et al. (2023). In stark contrast to their dielectric counterparts, plasmonic nanocavities offer the unique ability to confine light to extremely sub-wavelength scales and massively enhance local electromagnetic fields via their resonant surface plasmon modes, thereby constituting unique architectures for the exploration of light-matter interaction free from diffraction bounds Baumberg et al. (2019). Crucially, plasmonic resonators have ushered in an unprecedented era of room-temperature cQED at the nanoscale, facilitating access to the regime of strong coupling even at the profound limit of single molecular and quantum dot emitters under ambient conditions Chikkaraddy et al. (2016); Santhosh et al. (2016); Groß et al. (2018); Park et al. (2019), and opening up a compelling vista of potential applications in quantum nanotechnologies Xiong et al. (2021), extending from single-qubit coherent control Zhang et al. (2017) and ultrafast single-photon emission Hoang, Akselrod, and Mikkelsen (2016); Bogdanov et al. (2020); You et al. (2020) to near-field multipartite entanglement generation Bello et al. (2020); Bello, Kongsuwan, and Hess (2022) and quantum sensing Kongsuwan et al. (2019). In contrast to the electric dipole coupling between light and matter, cQED phenomena emerging from magnetic interactions have also been the subject of increasing attention. Leveraging collective microwave excitations of the spins in ferromagnetic and ferrimagnetic structures, the demonstration of such effects as photon-magnon and cavity-mediated magnon-qubit strong coupling Zhang et al. (2014); Hou and Liu (2019); Li et al. (2019); Neuman, Wang, and Narang (2020), have established cavity (nano-)magnonics as an attractive paradigm Rameshti et al. (2022), with interesting prospective applications in the engineering of synthetic gauge fields Gardin et al. (2024), cavity-mediated spin-photon interfaces at the nanoscale Chen et al. (2024), and the realization of quantum networks of entangled spins.
@@
Inevitably, a quantum description of lossy electromagnetic modes is essential towards the fundamental study and understanding of quantum nanophotonic systems, including the statistics of photon, plasmon or magnon emission, polaritonic effects and the role of quantum fluctuations, as well as to facilitate rigorous quantification of key figures of merit for emerging quantum nano-optical technologies. To date, most theoretical investigations of nanoscale cQED and spectroscopy (e.g., surface-enhanced and electron energy loss spectroscopies) have relied on an ad hoc quantization protocol, in which the bosonic cavity mode is directly modeled as a quantum harmonic oscillator without formal, field-theoretic justification, while decay effects are incorporated via the relevant dissipators in a quantum master equation treatment, themselves derived from a phenomenological coupling to a bath of continuum modes Neuman, Aizpurua, and Esteban (2020); Giulio, Kociak, and de Abajo (2019); Zhou et al. (2022). Recently however, significant progress was made with the introduction of a second quantization scheme for QNMs, enabling the construction of multiplasmon or multiphoton Fock states for arbitrary, three-dimensional (3D), dissipative and non-magnetic resonators Franke et al. (2019). This approach has been elaborated and applied to a range of problems, such as concerning the limits of (phenomenological) dissipative Jaynes-Cummings models in plasmonic cQED, the performance of near-field-driven single-photon sources Hughes et al. (2019), the description of nonlinear cQED effects in hybrid metal-dielectric cavity systems Franke et al. (2020a) and coupled lossy and amplifying resonators Franke, Ren, and Hughes (2022). An important feature of these works is the regularization of the QNMs based on a Dyson equation formalism and a dual expansion scheme for the dyadic Green’s function Ge et al. (2014), which utilizes the unregularized modes for positions inside the resonator and regularized modes that are functions of real frequency for positions outside. Whilst the Dyson equation approach appears to yield an accurate representation for the Green’s function outside the resonator volume and allows accurate calculation of the Purcell factor in this region for the studied systems Ge et al. (2014), numerical computation of the regularized fields appears computationally intensive, and may even become intractable for more complex photonic cavity systems. Moreover, the formal justification for such a mode regularization technique remains elusive Sauvan et al. (2022), and in particular, is based on an assumed completeness of the QNMs inside the resonator. Whilst this may apply to simple geometries, (i.e., compact resonators in a homogeneous background medium), it is not guaranteed for more complex geometries, such as resonators on a substrate or waveguide-coupled resonator systems Sauvan et al. (2022). Alternative quantization schemes for lossy and non-magnetic resonators have also emerged in the last few years, relying on a fitting of the spectral density of the electromagnetic environment Medina et al. (2021) based on emitter-centered modes Feist, Fernández-Domínguez, and García-Vidal (2020), or a transformation of the continuum of photonic eigenmodes into a discrete set of pseudomodes Yuen and Demetriadou (2024), where the latter offers the notable advantage of circumventing traditional reservoir approximations and can thereby capture non-Markovian effects in the dynamics of nanocavity-QE systems. Nevertheless, it remains desirable to formulate a unified and numerically feasible methodology for describing the lossy resonances of arbitrary, linear, magnetodielectric systems, directly addressing such crucial issues as mode divergence, the identification of canonical field variables in the presence of both dielectric and magnetic media, and the calculation of key figures of merit for quantum nano-optical device applications on rigorous foundations. It is worth mentioning here that the quantization of QNMs and their coupling to QEs is similar to the reaction coordinate mapping of a (strongly-coupled) spin-boson model with a structured bath, described by a spectral density peaked around a certain frequency Binder et al. (2018); Anto-Sztrikacs and Segal (2021). In this method the bosonic continuum is decomposed into a collective bosonic degree of freedom and a residual continuum for which the interaction can be treated within the Markov approximation.
@@
The paper is organized as follows. § II briefly reviews the classical electromagnetic QNM theory and PML regularization, as well as the auxiliary-field eigenvalue technique that yields a complete set of orthonormalized modes, alongside useful formulae for time-domain analysis, in the commonly treated case of non-dispersive magnetic permeabilities. In § III, we discuss the theory of mQED in arbitrary, linear, magnetodielectric media; here, we quantize the modes by incorporating PML regularization in the formalism and defining bosonic modal operators in order to construct the corresponding Fock space. Using these operators, we derive an effective Lindblad master equation in § IV to describe the dynamics of the quantized modes and their coupling to QEs. Finally, we close the paper in § V with some concluding remarks and perspectives.
@@ III Quantization of the regularized quasinormal modes
Similar to previous work Franke et al. (2019, 2020a); Franke, Ren, and Hughes (2022), our approach is predicated on the formalism of mQED, which offers a phenomenological Green’s function quantization scheme for electromagnetic fields in the presence of spatially inhomogeneous, lossy and dispersive media Gruner and Welsch (1996); Dung, Knöll, and Welsch (1998); Scheel, Knöll, and Welsch (1998); Dung, Knöll, and Welsch (2000); Raabe, Scheel, and Welsch (2007), later generalized for systems with gain Matloob et al. (1997); Raabe and Welsch (2008). However, in order to facilitate a rigorous application of the theory, a brief discussion concerning some important prerequisites is warranted. The complex permittivity $\varepsilon(\bm{r},\omega)=\varepsilon_{\mathrm{r}}(\bm{r},\omega)+\mathrm{i}\varepsilon_{\mathrm{i}}(\bm{r},\omega)$ and inverse permeability $\mu^{-1}(\bm{r},\omega)=\kappa(\bm{r},\omega)=\kappa_{\mathrm{r}}(\bm{r},\omega)+\mathrm{i}\kappa_{\mathrm{i}}(\bm{r},\omega)$ are required to satisfy the Kramers-Kronig relations (i.e., they are analytic in the closed, upper-half, complex $\omega$ plane and $\zeta(\bm{r},\omega)\rightarrow 1$ for $|\omega|\rightarrow\infty$ ($\zeta\in\{\varepsilon,\kappa\}$). This is not really a requirement of the quantum theory, but arises from the physical necessity of causality. Furthermore, they should be reciprocal, $\zeta(\bm{r},\omega)=\zeta^{\mathrm{T}}(\bm{r},\omega)$, and satisfy the reflection symmetry condition $\zeta_{ij}(\bm{r},-\omega^{*})=\zeta^{*}_{ij}(\bm{r},\omega)$, which guarantees their reality in the time domain. These properties ensure that the fundamental QED commutation relations between the electric and magnetic field operators are maintained. If the material parameters $\zeta(\bm{r},\omega)$ of the system are reciprocal, it readily follows from Eq. (10) that the PML-transformed parameters $\tilde{\zeta}(\bm{r},\omega)$ are as well. It should be noted that there have been generalizations of mQED for the cases of non-reciprocal materials and materials with magnetoelectric crossing effects Buhmann, Butcher, and Scheel (2012), which is, however, beyond the scope of this work.
@@ 1. Combination of electric/magnetic operators
$\bm{E}^{(+)}(\bm{r})=\sum_{n}\tilde{\bm{E}}^{\mathrm{sym}}_{n,\mathrm{L}}(\bm{r})\hat{a}_{n,\mathrm{L}}+\tilde{\bm{E}}^{\mathrm{sym}}_{n,\mathrm{G}}(\bm{r})\hat{a}^{\dagger}_{n,\mathrm{G}}\ ,$ (58)
@@ 2. Combination of gain/loss operators
$\bm{E}^{(+)}(\bm{r})=\sum_{n}\tilde{\bm{E}}^{\mathrm{sym}}_{n,\mathrm{e}}(\bm{r})\hat{a}_{n,\mathrm{e}}+\tilde{\bm{E}}^{\mathrm{sym}}_{n,\mathrm{m}}(\bm{r})\hat{a}_{n,\mathrm{m}}\ ,$ (60)

### v2

@@ Abstract
The unprecedented pace of evolution in nanoscale architectures for cavity quantum electrodynamics (cQED) has posed crucial challenges for theory, where the quantum dynamics arising from the non-perturbative dressing of matter by cavity electric and magnetic fields, as well as the fundamentally non-hermitian character of the system are to be treated without significant approximation. The lossy electromagnetic resonances of photonic, plasmonic or magnonic nanostructures are described as quasinormal modes (QNMs), whose properties and interactions with quantum emitters and spin qubits are central to the understanding of dissipative nano-optics and magnetodielectric cQED. Despite recent advancements toward a fully quantum framework for QNMs, a general and universally accepted approach to QNM quantization for arbitrary linear media remains elusive. In this work, we introduce a unified theoretical framework, based on macroscopic QED and complex coordinate transformations, that achieves QNM quantization for a wide class of spatially inhomogeneous, dissipative (with possible gain components) and dispersive, linear, magnetodielectric resonators. The complex coordinate transformations equivalently convert the radiative losses into non-radiative material dissipation, and via a suitable transformation that reflects all the losses of the resonator, we define creation and annihilation operators that allow the construction of modal Fock states for the joint excitations of field-dressed matter. By directly addressing the intricacies of modal loss in a fully quantum theory of magnetodielectric cQED, our approach enables the exploration of modern, quantum nano-optical experiments utilizing dielectric, plasmonic, magnetic or hybrid cQED architectures, and paves the way towards a rigorous assessment of room-temperature quantum nanophotonic technologies without recourse to ad hoc quantization schemes.
@@ I Introduction
Cavity-mediated control and enhancement of light-matter interaction has become a pervasive paradigm in modern quantum science and technology. Already decades in maturity, the use of micro- and nanoscale, optical, dielectric resonators to both harness and manipulate the photonic properties of semiconductor quantum dots has enabled the exploration of fundamental cavity quantum electrodynamics (cQED) in the solid state Vahala (2003); Khitrova et al. (2006); Kira and Koch (2011), as well as a host of device applications ranging from high-performance, non-classical light sources and spin-photon interfaces to ultrafast, all-optical switching for integrated photonic quantum computation and networks Hepp et al. (2019); Arakawa and Holmes (2020); Heindel et al. (2023). In stark contrast to their dielectric counterparts, plasmonic nanocavities offer the unique ability to confine light to extremely sub-wavelength scales and massively enhance local electromagnetic fields via their resonant surface plasmon modes, thereby constituting unique architectures for the exploration of light-matter interaction free from diffraction bounds Baumberg et al. (2019). Crucially, plasmonic resonators have ushered in an unprecedented era of room-temperature cQED at the nanoscale, facilitating access to the regime of strong coupling even at the profound limit of single molecular and quantum dot emitters under ambient conditions Chikkaraddy et al. (2016); Santhosh et al. (2016); Groß et al. (2018); Park et al. (2019), and opening up a compelling vista of potential applications in quantum nanotechnologies Xiong et al. (2021); Clarke and Hess (2024), extending from single-qubit coherent control Zhang et al. (2017) and ultrafast single-photon emission Hoang, Akselrod, and Mikkelsen (2016); Bogdanov et al. (2020); You et al. (2020) to near-field multipartite entanglement generation Bello et al. (2020); Bello, Kongsuwan, and Hess (2022) and quantum sensing Kongsuwan et al. (2019). In contrast to the electric dipole coupling between light and matter, cQED phenomena emerging from magnetic interactions have also been the subject of increasing attention. Leveraging collective microwave excitations of the spins in ferromagnetic and ferrimagnetic structures, the demonstration of such effects as photon-magnon and cavity-mediated magnon-qubit strong coupling Zhang et al. (2014); Hou and Liu (2019); Li et al. (2019); Neuman, Wang, and Narang (2020), have established cavity (nano-)magnonics as an attractive paradigm Rameshti et al. (2022), with interesting prospective applications in the engineering of synthetic gauge fields Gardin et al. (2024), cavity-mediated spin-photon interfaces at the nanoscale Chen et al. (2024), and the realization of quantum networks of entangled spins.
@@
Inevitably, a quantum description of lossy electromagnetic modes is essential towards the fundamental study and understanding of quantum nanophotonic systems, including the statistics of photon, plasmon or magnon emission, polaritonic effects and the role of quantum fluctuations, as well as to facilitate rigorous quantification of key figures of merit for emerging quantum nano-optical technologies. To date, most theoretical investigations of nanoscale cQED and spectroscopy (e.g., surface-enhanced and electron energy loss spectroscopies) have relied on an ad hoc quantization protocol, in which the bosonic cavity mode is directly modeled as a quantum harmonic oscillator without formal, field-theoretic justification, while decay effects are incorporated via the relevant dissipators in a quantum master equation treatment, themselves derived from a phenomenological coupling to a bath of continuum modes Neuman, Aizpurua, and Esteban (2020); Giulio, Kociak, and de Abajo (2019); Zhou et al. (2022). Recently however, significant progress was made with the introduction of a second quantization scheme for QNMs, enabling the construction of multiplasmon or multiphoton Fock states for arbitrary, three-dimensional (3D), dissipative and non-magnetic resonators Franke et al. (2019). This approach has been elaborated and applied to a range of problems, such as concerning the limits of (phenomenological) dissipative Jaynes-Cummings models in plasmonic cQED, the performance of near-field-driven single-photon sources Hughes et al. (2019), the description of nonlinear cQED effects in hybrid metal-dielectric cavity systems Franke et al. (2020a) and coupled lossy and amplifying resonators Franke, Ren, and Hughes (2022). An important feature of these works is the regularization of the QNMs based on a Dyson equation formalism and a dual expansion scheme for the dyadic Green’s function Ge et al. (2014), which utilizes the unregularized modes for positions inside the resonator and regularized modes that are functions of real frequency for positions outside. Whilst the Dyson equation approach appears to yield an accurate representation for the Green’s function outside the resonator volume and allows accurate calculation of the Purcell factor in this region for the studied systems Ge et al. (2014), numerical computation of the regularized fields appears computationally intensive, and may even become intractable for more complex photonic cavity systems. Moreover, the formal justification for such a mode regularization technique remains elusive Sauvan et al. (2022), and in particular, is based on an assumed completeness of the QNMs inside the resonator. Whilst this may apply to simple geometries, (i.e., compact resonators in a homogeneous background medium), it is not guaranteed for more complex geometries, such as resonators on a substrate or waveguide-coupled resonator systems Sauvan et al. (2022). Alternative quantization schemes for lossy and non-magnetic resonators have also emerged in the last few years, relying on a fitting of the spectral density of the electromagnetic environment Medina et al. (2021) based on emitter-centered modes Feist, Fernández-Domínguez, and García-Vidal (2020), or a transformation of the continuum of photonic eigenmodes into a discrete set of pseudomodes Yuen and Demetriadou (2024), where the latter offers the notable advantage of circumventing traditional reservoir approximations and can thereby capture non-Markovian effects in the dynamics of nanocavity-QE systems. Nevertheless, it remains desirable to formulate a unified and numerically feasible methodology for describing the lossy resonances of arbitrary, linear, magnetodielectric systems, directly addressing such crucial issues as mode divergence, the identification of canonical field variables in the presence of both dielectric and magnetic media, and the calculation of key figures of merit for quantum nano-optical device applications on rigorous foundations. It is worth mentioning here that the quantization of QNMs and their coupling to QEs is similar to the reaction coordinate mapping of a (strongly-coupled) spin-boson model with a structured bath, described by a spectral density peaked around a certain frequency Binder et al. (2018); Anto-Sztrikacs and Segal (2021). In this method, the bosonic continuum is decomposed into a collective bosonic degree of freedom and a residual continuum, for which the interaction can be treated within the Markov approximation.
@@
The paper is organized as follows. § II briefly reviews the classical electromagnetic QNM theory and PML regularization, as well as the auxiliary-field eigenvalue technique that yields a complete set of orthonormalized modes, alongside useful formulae for time-domain analysis, in the commonly treated case of non-dispersive magnetic permeabilities. In § III, we discuss the theory of mQED in arbitrary, linear, magnetodielectric media; here, we quantize the modes by incorporating PML regularization in the formalism and defining bosonic modal operators in order to construct the corresponding Fock space. Using these operators, we derive an effective Lindblad master equation in § IV to describe the dynamics of the quantized modes and their coupling to QEs. In § V, we demonstrate the theory by investigating the spontaneous decay of a QE in a simple, 1D, half-open cavity, comparing our predictions with the semi-analytical results of an exact mQED treatment. Finally, we close the paper in § VI with some concluding remarks and perspectives.
@@ III Quantization of the regularized quasinormal modes
Similar to previous work Franke et al. (2019, 2020a); Franke, Ren, and Hughes (2022), our approach is predicated on the formalism of mQED, which offers a phenomenological Green’s function quantization scheme for electromagnetic fields in the presence of spatially inhomogeneous, lossy and dispersive media Gruner and Welsch (1996); Dung, Knöll, and Welsch (1998); Scheel, Knöll, and Welsch (1998); Dung, Knöll, and Welsch (2000); Raabe, Scheel, and Welsch (2007), later generalized for systems with gain Matloob et al. (1997); Raabe and Welsch (2008). However, in order to facilitate a rigorous application of the theory, a brief discussion concerning some important prerequisites is warranted. The complex permittivity $\varepsilon(\bm{r},\omega)=\varepsilon_{\mathrm{r}}(\bm{r},\omega)+\mathrm{i}\varepsilon_{\mathrm{i}}(\bm{r},\omega)$ and inverse permeability $\mu^{-1}(\bm{r},\omega)=\kappa(\bm{r},\omega)=\kappa_{\mathrm{r}}(\bm{r},\omega)+\mathrm{i}\kappa_{\mathrm{i}}(\bm{r},\omega)$ are required to satisfy the Kramers-Kronig relations (i.e., they are analytic in the closed, upper-half, complex $\omega$ plane and $\zeta(\bm{r},\omega)\rightarrow 1$ for $|\omega|\rightarrow\infty$ ($\zeta\in\{\varepsilon,\kappa\}$). This is not really a requirement of the quantum theory, but arises from the physical necessity of causality. Furthermore, they should be reciprocal, $\zeta(\bm{r},\omega)=\zeta^{\mathrm{T}}(\bm{r},\omega)$, and satisfy the reflection symmetry condition $\zeta_{ij}(\bm{r},-\omega^{*})=\zeta^{*}_{ij}(\bm{r},\omega)$, which guarantees their reality in the time domain. These properties ensure that the fundamental QED commutation relations between the electric and magnetic field operators are maintained. If the material parameters $\zeta(\bm{r},\omega)$ of the system are reciprocal, it readily follows from Eq. (10) that the PML-transformed parameters $\tilde{\zeta}(\bm{r},\omega)$ are as well. It should be noted that there have been generalizations of mQED for the cases of non-reciprocal materials and materials with magnetoelectric crossing effects Buhmann, Butcher, and Scheel (2012), which are, however, beyond the scope of this work.
@@ 1. Combination of electric/magnetic operators
$\hat{\bm{E}}^{(+)}(\bm{r})=\sum_{n}\tilde{\bm{E}}^{\mathrm{sym}}_{n,\mathrm{L}}(\bm{r})\hat{a}_{n,\mathrm{L}}+\tilde{\bm{E}}^{\mathrm{sym}}_{n,\mathrm{G}}(\bm{r})\hat{a}^{\dagger}_{n,\mathrm{G}}\ ,$ (58)
@@ 2. Combination of gain/loss operators
$\hat{\bm{E}}^{(+)}(\bm{r})=\sum_{n}\tilde{\bm{E}}^{\mathrm{sym}}_{n,\mathrm{e}}(\bm{r})\hat{a}_{n,\mathrm{e}}+\tilde{\bm{E}}^{\mathrm{sym}}_{n,\mathrm{m}}(\bm{r})\hat{a}_{n,\mathrm{m}}\ ,$ (60)